# Do We Truly Need So Many Samples? Multi-LLM Repeated Sampling Efficiently Scale Test-Time Compute

### v1

@@ Abstract
This paper presents a simple, effective, and cost-efficient strategy to improve LLM performance by scaling test-time compute. Our strategy builds upon the repeated-sampling-then-voting framework, with a novel twist: incorporating multiple models, even weaker ones, to leverage their complementary strengths that potentially arise from diverse training data and paradigms. By using consistency as a signal, our strategy dynamically switches between models. Theoretical analysis highlights the efficiency and performance advantages of our strategy. Extensive experiments on six datasets demonstrate that our strategy not only outperforms self-consistency and state-of-the-art multi-agent debate approaches, but also significantly reduces inference costs. Additionally, ModelSwitch requires only a few comparable LLMs to achieve optimal performance and can be extended with verification methods, demonstrating the potential of leveraging multiple LLMs in the generation-verification paradigm.
@@ 1 Introduction
We conduct extensive evaluations on six datasets that cover a wide range of knowledge, reasoning, comprehension, and domain-specific challenges. ModelSwitch, when leveraging Gemini 1.5 Flash and GPT-4o mini, significantly surpasses the self-consistency [17] performance of each respective LLM. Moreover, it even outperforms more advanced LLMs such as GPT-4o and Gemini 1.5 Pro—achievements that self-consistency cannot match. At the same time, ModelSwitch demonstrates superior sample efficiency, reducing the average sampling count per query by 34% on six datasets, showcasing its ability to save computational costs while maintaining high performance. Furthermore, ModelSwitch achieves state-of-the-art performance on five datasets, outperforming five other multi-agent or multi-LLM methods, and secures the second-best result on the remaining dataset. Notably, ModelSwitch achieves an impressive 63.2% accuracy on the MMLU-Pro [20] dataset, surpassing the best single LLM by 10.2 points and significantly exceeding other methods, including MAD [21] (45%), AgentVerse [22] (38%), ChatEval [23] (43%), MAD (multi-LLM version) (50.2%), and MOA [16] (50.8%). In additional experiments, we demonstrate that ModelSwitch requires only a small number of LLMs with comparable performance to achieve optimal results and is relatively robust to the order of models. We also show that ModelSwitch can be integrated with verification methods, such as reward models, further enhancing the performance.
We conduct a formal analysis to better understand the improvements brought by ModelSwitch, compared to self-consistency. For a given query, we derive a sufficient condition and a necessary condition that ModelSwitch (two models) can surpass the performance of single model. This sufficient condition may still be satisfied even when neither of the two models can produce a completely consistent answer lists. This reveals the source of the superiority of ModelSwitch compared to a single model. Furthermore, we provide a theoretical bound on the efficiency gains achieved by ModelSwitch, offering deeper insights into its advantages. In particular, if the probability of each model producing a completely consistent answer list is greater than some constant $c>0$ and each model is sampled the same number of times, the expected number of samplings can be decreased by a factor $\frac{1}{n}\frac{1-c}{c}$, where $n$ is the number of models.
@@ 3 ModelSwitch: Harnessing Consistency in Multi-LLM Generation
Following the sampling-then-voting strategy, the most common method is majority voting, where the most frequently occurring answer among all LLMs is selected as the final answer. However, when a large quantity of models are involved in voting, performance differences often exist among them, so equally aggregating the answers from all models is not an optimal approach. Therefore, we design a weighted voting algorithm with internal weights $\text{W}_{\alpha}$ and external weights $\text{W}_{\beta}$. Each answer will be multiplied by these two weights when voting.
@@
[…]
Consider that we have 3 model answer sets denoted as $\text{M}_{1}$, $\text{M}_{2}$, $\text{M}_{3}$. $\text{M}_{1}=\{\text{A}*3,\text{B}*2\}$, $\text{M}_{2}=\{\text{B}*3,\text{A}*2\}$ , $\text{M}_{3}=\{\text{C}*5\}$ According to the majority voting algorithm, answers A, B, and C have the same score. However, since $\text{M}_{3}$ has higher confidence, the score of C becomes higher after incorporating internal weights $\text{W}_{\alpha}$, allowing it to stand out. Our goal is to ensure that the answers of the more confident models receive higher scores.
External weights $\text{W}_{\beta}$ assess prior performance, ensuring models with significantly different performances have corresponding importance in voting. For example, we can assign $\text{W}_{\beta}=$ 2 or 1.5 to much stronger models, while giving $\text{W}_{\beta}=$ 1 to weaker models. Specifically, when there are only two models, we set $\text{W}_{\beta}=$ uniformly to 1.
@@ 4.1 Experimental Setup
We mainly evaluate our method on the following six datasets that cover a wide range of knowledge, reasoning, comprehension, and domain-specific challenges, providing a robust evaluation for our method: GSM8K [8], MATH [18], MathBench [19], MGSM [24], DATE [25], and MMLU-Pro [20]. Details of the six datasets are provided in Section A.1.
@@
We primarily consider three lightweight, closed-source LLMs: GPT-4o mini [26], Gemini 1.5 Flash [27], and Claude 3 Haiku [28]. In Section 4.2, we compare ModelSwitch to single-LLM repeated sampling methods, using GPT-4o mini and Gemini 1.5 Flash, while also including more advanced LLMs, GPT-4o and Gemini 1.5 Pro, as additional baselines. Subsequently, in Section 4.3, we use all three lightweight LLMs to evaluate ModelSwitch against other multi-agent debate methods. In Section 4.4, we investigate the impact of the number of LLMs by combining the three aforementioned lightweight closed-source LLMs with three open-source LLMs—Llama-3.1-8B-Instruct [29], Gemma-2-9B-It [30], and Qwen2.5-7B-Instruct [9]—to assess how performance scales with model diversity. Finally, in Section 4.5, we explore whether ModelSwitch can be enhanced by integrating stronger validation methods, using Qwen2.5-MATH-RM-72B. For details on all hyperparameter settings, see Section A.2.
@@ 4.3 ModelSwitch Better Leverages Multi-LLM or Multi-Agent Synergies than Debate Methods
Figure 5: Performance comparison of different multi-agent debate systems under the same budget of 15 samples relative to the accuracy of the best single LLM with one sample. ModelSwitch achieves the best results on five datasets and the second-best result on DATE.
@@
As shown in Figure 5, our method achieves state-of-the-art performance on five datasets, with a significant lead on MMLU-Pro. For example, MAD (45%), AgentVerse (38%), ChatEval (43%), MLD (50.2%) and MOA (50.8%) all perform worse than the best single LLM (53%) sampling once on MMLU-Pro. This performance gap highlights a key limitation of these systems: while they rely on interactions between multiple agents (LLMs), it is difficult to ensure that such interactions consistently guide the answers in the correct direction. Challenges like error propagation often arise during these interactions [15, 31], which are especially hard to mitigate on a complex dataset like MMLU-Pro. In contrast, ModelSwitch achieves a remarkable 63.2% accuracy, representing an impressive 10.2-point increase over the best single LLM. This demonstrates its ability to effectively address these issues and deliver superior performance. This also highlights the potential of exploring how to effectively promote collaboration among different LLMs as a promising research direction.
@@ 5 Understanding Why ModelSwitch Outperforms Self-Consistency
Consider a set of $n$ models, denoted as $\{M_{1},M_{2},\ldots,M_{n}\}$. For a given query and a total sampling budget $K$, each model is sampled sequentially up to $\frac{K}{n}$ times, such that the theoretical maximum number of samples is $K$. Each model $M_{i}$ has a probability $P_{i}$ of producing a completely consistent answer list. If a model produces such a consistent answer list, subsequent models are pruned, and no further sampling is performed.
@@
$\mathbb{E}[N_{\text{call}}]=\frac{K}{n}\sum_{k=1}^{n}\prod_{i=1}^{k}(1-P_{i}).$ (2)
If the probability of each model producing a completely consistent answer list is greater than some constant $c>0$, the expected number of samplings can be bounded by $\frac{K}{n}\frac{1-c}{c}$, which reduces the number of samplings by a factor $\frac{1-c}{nc}$.
@@ 7 Conclusion
Appendix A Detailed Experimental Setup
A.1 Benchmarks
The details of the six datasets we used are as follows:
@@ A.1 Benchmarks
A.2 Hyperparameter Settings
@@
The settings for external weights $\text{W}_{\beta}$ in Sections 4.3 and 4.4 are shown in Tables 1 and 2, respectively.
@@
GPT-4o mini 1 2 1 1 1.5 1
Gemini 1.5 Flash 1 2 1 1 1.5 1.5
@@
A.3 Multi-Agent Debate Systems Setup
@@ C.4 Validation of Theoretical Formulation
By choosing an $\epsilon$ that satisfies Equation 5, and combining Equations 7 and 9, we have completed the proof of this theorem.

### v4

# Do We Truly Need So Many Samples? Multi-LLM Repeated Sampling Efficiently Scales Test-Time Compute

@@ Abstract
This paper presents a simple, effective, and cost-efficient strategy to improve LLM performance by scaling test-time compute. Our strategy builds upon the repeated-sampling-then-voting framework, with a novel twist: incorporating multiple models, even weaker ones, to leverage their complementary strengths that potentially arise from diverse training data and paradigms. By using consistency as a signal, our strategy dynamically switches between models. Theoretical analysis highlights the efficiency and performance advantages of our strategy. Extensive experiments on seven datasets demonstrate that our strategy not only outperforms self-consistency and state-of-the-art multi-agent debate approaches, but also significantly reduces inference costs. Additionally, ModelSwitch requires only a few comparable LLMs to achieve optimal performance and can be extended with verification methods, demonstrating the potential of leveraging multiple LLMs in the generation-verification paradigm.
@@ 1 Introduction
We conduct extensive evaluations on seven datasets that cover a wide range of knowledge, reasoning, comprehension, and domain-specific challenges. ModelSwitch, when leveraging Gemini 1.5 Flash and GPT-4o mini, significantly surpasses the self-consistency [17] performance of each respective LLM. Moreover, it even outperforms more advanced LLMs such as GPT-4o and Gemini 1.5 Pro—achievements that self-consistency cannot match. At the same time, ModelSwitch demonstrates superior sample efficiency, reducing the average sampling count per query by 34% on six datasets, showcasing its ability to save computational costs while maintaining high performance. Furthermore, ModelSwitch achieves state-of-the-art performance on four datasets, outperforming five other multi-agent or multi-LLM methods, and secures the second-best result on the remaining two datasets. Notably, ModelSwitch achieves an impressive 63.2% accuracy on the MMLU-Pro [20] dataset, surpassing the best single LLM by 10.2 points and significantly exceeding other methods, including MAD [21] (47.6%), AgentVerse [22] (44.8%), ChatEval [23] (44%), MAD (multi-LLM version) (44%), and MOA [16] (52.6%). In additional experiments, we demonstrate that ModelSwitch requires only a small number of LLMs with comparable performance to achieve optimal results and is relatively robust to the order of models. We also show that ModelSwitch can be integrated with verification methods, such as reward models, further enhancing the performance.
We conduct a formal analysis to better understand the improvements brought by ModelSwitch, compared to self-consistency. For a given query, we derive a sufficient condition and a necessary condition that ModelSwitch (two models) can surpass the performance of single model. This sufficient condition may still be satisfied even when neither of the two models can produce a completely consistent answer lists. This reveals the source of the superiority of ModelSwitch compared to a single model. Furthermore, we provide a theoretical bound on the efficiency gains achieved by ModelSwitch, offering deeper insights into its advantages. In particular, if the probability of each model producing a completely consistent answer list is greater than some constant $c>0$ and each model is sampled the same number of times, the expected number of samplings can be decreased by a factor $\frac{1}{n\times c}$, where $n$ is the number of models.
@@ 3 ModelSwitch: Harnessing Consistency in Multi-LLM Generation
To account for performance differences among models, we use a weighted voting algorithm with internal ($\text{W}_{\alpha}$) and external ($\text{W}_{\beta}$) weights to aggregate answers $A$ from all models $M$.
$MS(q)=\operatorname*{argmax}_{a\in A}\sum_{i=1}^{M}\text{W}_{\alpha}\times\text{W}_{\beta}\times\text{count}(a,A_{i})$ (1)
Internal Weights: Measure a model’s confidence based on answer consistency. Calculated via entropy and normalized to $[\text{bias},1]$, where $\text{bias}=\frac{1}{\text{len(samples)}}$. Higher consistency yields a higher weight. Example: For answer sets $\text{M}_{1}=\{\text{A}\times 3,\text{B}\times 2\}$, $\text{M}_{2}=\{\text{B}\times 3,\text{A}\times 2\}$, $\text{M}_{3}=\{\text{C}\times 5\}$, $\text{M}_{3}$’s uniformity gives C a higher score.
@@
External Weights: Assess prior performance, ensuring models with significantly different performances have corresponding importance. Specific $\text{W}_{\beta}$ settings can be found in Section A.1.
@@ 4.1 Experimental Setup
We mainly evaluate our method on the following seven datasets that cover a wide range of knowledge, reasoning, comprehension, and domain-specific challenges, providing a robust evaluation for our method: GSM8K [8], MATH [18], MathBench [19], MGSM [24], DATE [25], MMLU-Pro [20] and AIME24. Details of the seven datasets are provided in Section A.1.
@@
We primarily consider three lightweight, closed-source LLMs: GPT-4o mini [26], Gemini 1.5 Flash [27], and Claude 3 Haiku [28]. In Section 4.2, we compare ModelSwitch to single-LLM repeated sampling methods, using GPT-4o mini and Gemini 1.5 Flash, while also including more advanced LLMs, GPT-4o and Gemini 1.5 Pro, as additional baselines. Subsequently, in Section 4.3, we use all three lightweight LLMs to evaluate ModelSwitch against other multi-agent debate methods. In Section 4.4, we investigate the impact of the number of LLMs by combining the three aforementioned lightweight closed-source LLMs with three open-source LLMs—Llama-3.1-8B-Instruct [29], Gemma-2-9B-It [30], and Qwen2.5-7B-Instruct [9]—to assess how performance scales with model diversity. Finally, in Section 4.5, we explore whether ModelSwitch can be enhanced by integrating stronger validation methods, using Qwen2.5-MATH-RM-72B. Section A.2 explores efficacy of ModelSwitch using reasoning LLMs Llama-3.1-Nemotron-Nano-8B-v1 [31] and DeepSeek-R1-Distill-Qwen-7B [32].
@@ 4.3 ModelSwitch Better Leverages Multi-LLM or Multi-Agent Synergies than Debate Methods
Figure 5: Performance comparison of different multi-agent debate systems under the same budget of 15 samples relative to the accuracy of the best single LLM with one sample. ModelSwitch achieves the best results on four datasets and the second-best result on MATH and DATE.
@@
As shown in Figure 5, our method achieves state-of-the-art performance on four datasets, with a significant lead on MMLU-Pro. For example, MAD [21] (47.6%), AgentVerse [22] (44.8%), ChatEval [23] (44%), MAD (multi-LLM version) (44%), and MOA [16] (52.6%) all perform worse than the best single LLM (53%) sampling once on MMLU-Pro. This performance gap highlights a key limitation of these systems: while they rely on interactions between multiple agents (LLMs), it is difficult to ensure that such interactions consistently guide the answers in the correct direction. Challenges like error propagation often arise during these interactions [15, 33], which are especially hard to mitigate on a complex dataset like MMLU-Pro. In contrast, ModelSwitch achieves a remarkable 63.2% accuracy, representing an impressive 10.2-point increase over the best single LLM. This demonstrates its ability to effectively address these issues and deliver superior performance. This also highlights the potential of exploring how to effectively promote collaboration among different LLMs as a promising research direction.
@@ 5 Understanding Why ModelSwitch Outperforms Self-Consistency
Consider a set of $n$ models, denoted as $\{M_{1},M_{2},\ldots,M_{n}\}$. For a given query and a total sampling budget $K$, each model is sampled sequentially up to $\frac{K}{n}$ times, such that the theoretical maximum number of samples is $K$. Each model $M_{i}$ has a probability $P_{i}$ of producing a completely consistent answer list, where $P_{n}=1$ indicates that if the first $n-1$ models do not produce a consistent answer, the $n$-th model must be sampled.. If a model produces such a consistent answer list, subsequent models are pruned, and no further sampling is performed.
@@
$\mathbb{E}[N_{\text{call}}]=\frac{K}{n}\sum_{i=1}^{n}\left(\prod_{j=1}^{i-1}(1-P_{j})\right)P_{i}$ (3)
If the probability of each model producing a completely consistent answer list is greater than some constant $c>0$, the expected number of samplings can be bounded by $\frac{K}{n\times c}$, which reduces the number of samplings by a factor $\frac{1}{n\times c}$.
@@ 7 Conclusion
Appendix A More Experimental Results
A.1 Experimental Setup
Benchmarks
The details of the seven datasets we used are as follows:
@@ A.1 Experimental Setup
  • AIME24: A prestigious high school mathematics competition dataset contains 30 challenging mathematical problems.
Hyperparameter Settings
@@
When there are only two models, we always set external weights $\text{W}_{\beta}=$ uniformly to 1. The settings for $\text{W}_{\beta}$ in Sections 4.3 and 4.4 are shown in Tables 1 and 2, respectively.
@@
GPT-4o mini 1 2 1 2 1.5 1
Gemini 1.5 Flash 1 2 1 2 1.5 1.5
@@
Multi-Agent Debate Systems Setup
@@
Compute Resources
All open-source LLMs in this paper were run on a cluster running Ubuntu 22.04, equipped with 1,600GB of memory and 8 NVIDIA A100 GPUs, each with 80GB of VRAM. While this setup provides ample resources, it is not strictly necessary to run these LLMs. For smaller LLMs (not more than 10 billion parameters), full precision inference can be performed on systems with as little as two NVIDIA RTX 4090 GPUs (24GB VRAM each). For the 70B LLMs, a minimum of four NVIDIA RTX 4090 GPUs or two NVIDIA A100 GPUs (80GB VRAM each) is required.
A.2 ModelSwitch using Reasoning LLMs on AIME24
In this section, we explore whether ModelSwitch is also applicable to reasoning LLMs. We still compare ModelSwitch with self-consistency. To ensure fair comparison, we evaluate self-consistency for each LLM (Llama-3.1-Nemotron-Nano-8B-v1 and DeepSeek-R1-Distill-Qwen-7B) and ModelSwitch using both, under the same sampling budget. The results are shown in Figure 8.
Figure 8: Performance comparison of ModelSwitch (MS) and self-consistency (SC) [17] on AIME24. MS switches between Llama-3.1-Nemotron-Nano-8B-v1 and DeepSeek-R1-Distill-Qwen-7B. The curves show SC performance for individual LLMs. ModelSwitch achieves equal or better performance with fewer samples at budget 8, 12 and 16.
When the sampling budget is 8 and 12, ModelSwitch achieves significant improvements over the self-consistency of the best-performing LLM (66.7% → 73.3%, 70% → 80%). When the sampling budget is 16, ModelSwitch achieves the same accuracy (76.7%) as the self-consistency of the best-performing LLM. The actual average sampling counts of ModelSwitch under these three budgets is 7.1, 10.8, and 14.7, respectively, saving an average of 9% in samples (due to the difficulty of the questions and the low probability of models producing completely consistent answers, the number of samples saved is relatively limited). Overall, ModelSwitch remains an effective method for reasoning LLMs—achieving equal or better performance with fewer samples.
A.3 Experimental Cost
To better understand ModelSwitch’s reduction in computational cost, we visually present the experimental expenses. The API call prices for GPT-4o mini and Gemini 1.5 Flash are identical: \$0.15 per 1M tokens for prompts and \$0.6 per 1M tokens for completions. Since the prompt (input) part is exactly the same, we only account for the cost of completions (output). Table 3 shows the actual costs for self-consistency using GPT-4o mini and Gemini 1.5 Flash, as well as for ModelSwitch using both, based on Figure 4 with a sampling budget of 16. Ensuring that ModelSwitch outperforms self-consistency, we also significantly reduce costs, consistent with the reduction in the number of samplings.
Price (\$) GSM8K MATH MathBench MGSM DATE MMLU-Pro
GPT-4o mini (SC) 2.1* 0.6 0.3* 1.8 0.5* 0.9
Gemini 1.5 Flash (SC) 1.4 0.6* 0.3 1.4* 0.3 0.7*
ModelSwitch $1.1_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 48\%}}$ $0.4_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 33\%}}$ $0.2_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 33\%}}$ $0.9_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 36\%}}$ $0.3_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 40\%}}$ $0.6_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 15\%}}$
Table 3: Cost comparison between ModelSwitch and self-consistency (SC). We use * to indicate the best-performing LLM on each dataset, and $\downarrow$ to denote the cost reduction of ModelSwitch compared to self-consistency using the best-performing LLM. Across six datasets, we achieved significant cost reductions while ensuring performance improvements.
A.4 Ablation Study
In this section, we investigate the improvements brought by the weighted voting algorithm in aggregating answers. As shown in Table 4, we conduct ablation study on the MATH, MGSM, DATE, and MMLU-Pro datasets, where both internal and external weights were used for voting. It is very clear that both internal weights and external weights are effective in voting.
Method MATH MGSM DATE MMLU-Pro
ModelSwitch 78.6 90.6 77.8 63.2
      w/o internal weights $78.4_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.2}}$ $90.5_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.1}}$ $77.0_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.8}}$ $62.0_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 1.2}}$
      w/o external weights $77.8_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.8}}$ $90.3_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.3}}$ $76.4_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 1.4}}$ $61.4_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 1.8}}$
      w/o internal and external weights $77.0_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 1.6}}$ $90.2_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 0.4}}$ $76.4_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 1.4}}$ $61.2_{{\color[rgb]{0.1953125,0.72265625,0.59375}\downarrow 2.0}}$
Table 4: Ablation study of the weighted voting algorithm.
@@ C.4 Validation of Theoretical Formulation
By choosing an $\epsilon$ that satisfies Equation 6, and combining Equations 8 and 10, we have completed the proof of this theorem.